# Emergence of Chimera states in an active fluid medium in the presence of chemical species with logistic growth

### v1

@@ I Introduction
Pattern formation is a ubiquitous phenomenon in natural systems. In particular, it is an integral part of the development of biological systems. The classical framework for modelling and understanding pattern formation is that of reaction-diffusion systems [1, 2, 3, 4, 5]. The emergent patterns observed in this broad class of systems include time-independent and time-dependent oscillatory patterns [1], symmetry-breaking instabilities [6], traveling waves [7], spirals [8] and jumping oscillations [9]. Importantly, many of these patterns have been verified in experiments as well [10, 11].
@@
However, in most biological systems, active transport and forces play a critical role in pattern formation. A prototypical approach to understanding pattern formation in biological systems, therefore, integrates the contributions of mechanical and chemical effects [12], for instance actin networks showing non-equilibrium dynamics by force generation through myosin motor activity [13, 14, 15] and pattern formation in active fluid medium in the presence of diffusing chemical species that are advected by self-generated flows produced by concentration-dependent active stress gradients [16, 13]. Specifically, the actomyosin cortex, which lies just beneath the cell membrane, consists of actin filaments and myosin motor proteins which crosslink the actin. This generates mechanical forces, giving rise to an active stress component in a thin mesoscopic layer of the actomyosin cortex. The actomyosin cortex is treated as an active fluid at developmental time scales. The active stress can now be considered to be a function of the concentrations of regulatory chemical species to complete the mechanochemical integration, leading to pattern formation even in the absence of chemical reactions [16]. When extended to two chemical species which undergo advection-diffusion and regulate the active stress, this framework gives rise to pulsatory patterns [13]. Spontaneously emerging localized states in the active fluid medium have led to the understanding of localized cellular patterns [17]. It has been shown that localized states can emerge spontaneously (analogous to isolated clusters of actin and signalling molecules in cancer cells [18]) if the assembly of active matter is regulated by the presence of chemical species that are advected with flows resulting from active stress gradients.
@@
In another research direction, in the context of dynamical systems in general, a class of patterns that have generated intense research interest are chimera states. Broadly, a chimera state is one where a system spontaneously breaks the underlying symmetry and splits into co-existing groups that have very different dynamical features [19, 20, 21, 22]. This fascinating phenomenon has been observed in a variety of systems, ranging from Josephson junction arrays [23] to uni-hemispheric sleep in certain animals [24], as well as in continuous media models [25, 26]. Importantly, these chimera states have also been observed experimentally in optical analogs of coupled map lattices [27], Belousov-Zhabotinsky chemical oscillator systems [28], two populations of mechanical metronomes [29], electronic circuit systems [30, 31], neurodynamics [32] and Liquid Crystal Light Valve experiments with optical feedback [33]. So it is of immense interest to uncover chimera patterns in different classes of systems in order to gauge the generality of this interesting spatiotemporal pattern.
@@
Motivated by these broad ideas, in this paper, we show the emergence of chimera states in an active fluid with two chemical species regulating the active stress as in Ref. [13], with one species showing a logistic growth. A logistic growth term has been used extensively in many mathematical models of biological systems and provides a very general description of population growth. In particular, a logistic growth model with long-range interactions serves as a generic minimal model for competition for common resources and pattern formation in excited media [34]. In the chemotaxis model, the logistic term is often used to model the growth of cell density [35]. In autochemotactic pattern formation, it is used to model the death and reproduction of self-propelled bacteria  [36]. Further, the logistic functional form has been used in an experimentally motivated model of the coupling between reaction-diffusion and active matter [37], as well as in gene expression dynamics [38, 39]. Polymerization in actin gels and solutions is also expressed by a reactive logistic term, useful to describe many characteristic states of actin-wave formation: spots, spirals, and travelling waves [40, 41].
@@ II Mathematical Model
Thus, both species undergo advection and diffusion, with the diffusive component determined by the diffusion coefficient $D$ and the advective component given by the bulk fluid flow velocity $v$. The relative diffusion coefficient of the two species is determined by the parameter $\alpha>0$. The nonlinear logistic growth term, with strength $r$, is a generic reaction term that has the capacity to destabilize a steady state with zero concentration and also saturates at a finite carrying capacity $K$. For the active fluid at low Reynolds numbers, the inertial terms can be neglected, and the force balance equation gives
@@
where $\beta$ is an asymmetry parameter, and $A_{S}$ and $I_{S}$ represent the saturation values of the concentrations of the two chemical species [13]. For $\beta<-1$, $A$ down-regulate and $I$ up-regulates active stress; for $-1\leq\beta\leq 1$ both species up-regulates stress and for $\beta>1$, $A$ up-regulates while $I$ down-regulates active stress.
@@
$\partial_{t}A=-\partial_{x}(vA)+D\partial^{2}_{x}A+RA(1-A)$ (7)
@@
III Stability Analysis
Figure 1: (Right) Phase diagram of the different dynamical patterns that emerge, in the parameter space of the Péclet number $Pe$ and the scaled nonlinear growth parameter $R$, obtained from Linear Stability Analysis. Here the ratio of diffusion constants $\alpha=0.1$ and the asymmetry parameter $\beta=3$. The brown colour represents the stable homogeneous steady state, pale blue represents oscillatory instability and yellow represents stationary instability. The black curve represents the approximate linear boundary between the stable homogeneous steady state and the regime of oscillatory instability, given by Eqn. 13. Left panels (a),(b) and (c) display the dispersion curves corresponding to $Pe<Pe_{c}$, $Pe=Pe_{c}$ and $Pe>Pe_{c}$ respectively, where $Pe_{c}=11.2$ is the critical Péclet number. The blue curve represents the real part of the leading eigenvalue, the red curve represents the imaginary part of the leading eigenvalue, and the black dotted line shows $\lambda=0$ for reference.
We consider perturbations to the homogeneous steady state $(A_{0},I_{0})$ of the form $A=A_{0}+A_{p}(x,t)$, $I=I_{0}+I_{p}(x,t)$ and $v=v_{0}+v_{p}(x,t)$, where, $A_{p}(x,t)=\delta A_{0}e^{\lambda t+ikx}$, $I_{p}(x,t)=\delta I_{0}e^{\lambda t+ikx}$, $v_{p}(x,t)=\delta v_{0}e^{\lambda t+ikx}$. The important parameters here are (1) the Péclet number, which gives the ratio of the advective time scale and the diffusive time scale, (2) the parameter $R$, which reflects the strength of nonlinearity, and (3) $\beta$, which reflects the asymmetry in the two chemical species and influences evolution through the function $f$. Linear stability analysis of Eqs. (7)-(9), yields the following Jacobian:
@@ III Stability Analysis
where $f_{A},f_{I}$ are the partial derivatives of $f$ with respect to $A$ and $I$ respectively, evaluated at the homogeneous steady state. In what follows we choose $(A_{S},I_{S})=(3A_{0},3I_{0})$. For each value of wave number $k=\frac{2\pi n}{L}$, $n=0,1,\dots$ (considering periodic boundary conditions), we obtain two eigenvalues, which determine the stability and nature of the emergent dynamical pattern. The trace and the determinant of the Jacobian are given as:
@@
Now, for a homogeneous steady state to be stable under perturbation, the disturbances from the homogeneous profile must damp exponentially rapidly in time. So the real parts of the eigenvalues for all $k$ should be less than zero, implying that $Tr(J)<0$ and $\Delta(J)>0$. In the parameter regime where the uniform steady state is unstable, the real part of at least one of the eigenvalues is positive, leading to the formation of diverse patterns. We denote the regime where the leading eigenvalues may not be imaginary as a stationary instability, in analogy with the terminology used in the Turing model. If the eigenvlaues are imaginary for some values of $k$, an oscillatory instability will emerge.
Figure 2: Phase diagram of the different dynamical patterns that emerge, in the parameter space of the Péclet number $Pe$ and asymmetry parameter $\beta$, obtained from Linear Stability Analysis, for $\alpha$=$0.1$ and scaled nonlinear growth parameter $R$ equal to (a) $0.5$ and (b) $2$.
@@
In Fig. 1, we display the phase diagram in the parameter space of the Péclet number $Pe$ and the scaled nonlinear growth parameter $R$. The phase diagram is generated considering the variation of the eigenvalues with respect to wave number $k$, and noting the sign of the real part of the leading eigenvalue. The three distinct dynamical phases, namely the stable homogeneous steady state, the regime of oscillatory instability and the regime of stationary instability, are marked with different colors.
@@
It is apparent from the phase diagram that as the Péclet number increases beyond a critical value, the homogeneous steady state gets destabilized giving way to an oscillatory instability. A set of dispersion curves (a-c) are also displayed along-side the phase diagram. These show the behaviour of the real and imaginary parts of the leading eigenvalue when the Péclet number crosses the critical value, thus illustrating the transition from the stable homogeneous steady state to oscillatory instability. On further increase of the Péclet number, the oscillatory instability changes to a stationary instability.
@@
In contrast, interestingly, the phase diagram also shows that the parameter space occupied by the homogeneous steady state is enlarged under increasing $R$, i.e. larger nonlinear growth yields enhanced homogeneity, indicating the counter-intuitive stabilizing effect of logistic growth on the dynamics. Further, increasing the magnitude of the nonlinear growth term also increases the parameter region supporting oscillatory states. It is also distinctly evident from Figure 1 that the boundaries between the different dynamical regimes have a linear dependence on the scaled nonlinearity $R$. The slope of the linear rise of the boundary curve separating the homogeneous steady state and the oscillatory state is gentler than that of the boundary between the stationary instability and oscillatory instability. This implies that the oscillatory state increases most significantly with increasing nonlinear growth.
@@
Specifically, our analysis also yields the boundary curve between the stable homogeneous steady state and the oscillatory instability. Along this boundary curve, the condition $Tr(J)=0$ should be satisfied. Since $k_{max}$ is close to $1$ over a large range of Péclet numbers, we can get an approximate expression for the critical Péclet number $Pe_{c}$ beyond which instability sets in. Considering the values of $\beta$, $\alpha$, $A_{0}$, $I_{0}$, $A_{S}$ and $I_{S}$ given in the phase diagram, this yields:
@@
This linear relation is indicated by a line in the phase diagram displayed in Fig. 1, and it is clear that this approximate expression fits results from stability analysis over a large large of $R$. Since the critical value of the Péclet number increases almost linearly with the scaled nonlinear growth parameter $R$, it implies that the range of $Pe$ over which the homogeneous steady state is stable is directly proportional to $R$, with larger stable ranges emerging for larger $R$.
@@
Figure 3: Dispersion curves obtained from Linear Stability Analysis, for varying Péclet number $Pe$ and asymmetry parameter $\beta$. Here $\alpha=0.1$ and nonlinearity parameter $R=1$. The first row of panels (a), (b) and (c), corresponding to $Pe<Pe_{c}$, $Pe=Pe_{c}$ and $Pe>Pe_{c}$, show the transition from the stable homogeneous steady state to the oscillatory instability, as $Pe$ increases from $9$ to $15$, for $\beta=1.2$. Panels (d), (e) and (f) show the transition from the stable homogeneous steady state to stationary instability as $\beta$ decreases from $1.1$ to $0.7$, for fixed Péclet number $Pe=6$. In all panels the blue curve represents the real part of the leading eigenvalue, the red curve represents the imaginary part of the leading eigenvalue, and the black dotted line shows $\lambda=0$ for reference.
Figure 4: (Left) Kymographs displaying the evolution of concentration $A$ in time (vertical axis) and space (horizontal axis), for parameter values (a) $Pe=13$, $\beta=1.5$, $R=1$, $\alpha=0.1$, and (b) $Pe=48$, $\beta=3$, $R=1$, $\alpha=0.1$. The colors in the heat-map of the concentration $A$ is given by the color-bar that indicates increasing concentrations from $A=0$ (black) to $A=2$ (white). (Right) Dynamical attractors in the phase space of $A-I$ concentrations at a typical location in space: (c) Limit Cycle corresponding to the space-time evolution shown in (a), and (d) Chaotic attractor corresponding to the space-time evolution shown in (b).
Figure 5: Emergence of an illustrative sequence of space-time patterns, ranging from transient Chimera states to irregular arrhythmic spatio-temporal patterns, in a system with $R=1$, $\alpha=0.1$, $\beta=3$, for increasing Péclet numbers: (a) $Pe=27$,(b) $Pe=35$, (c) $Pe=39$,(d) $Pe=48$. Here the heat-map of the concentration $A$ is displayed as a function of time (vertical axis) and space (horizontal axis), with the color bar indicating increasing concentrations from $A=0$ (black) to $A=3$ (white).
@@
We also observe pattern formation is crucially dependent on the asymmetry parameter $\beta$. This is clearly evident from Figure 2, where we display phase diagrams in the parameter space of the Péclet number $Pe$ and asymmetry $\beta$ for different values of the scaled logistic growth parameter $R$. When $\beta$ is low ($\beta<1$), no regime of oscillatory instability exists for any value of the Péclet number, while for $\beta$ greater than $1$ all three dynamical regimes are possible. Two sets of dispersion curves are also displayed in Fig. 3, showing the behaviour of the real and imaginary parts of the leading eigenvalues when $Pe$ and $\beta$ cross critical values, underlying different classes of dynamical transitions. In particular, these dispersion curves illustrate the transitions from the stable homogeneous steady state to oscillatory instability on increasing $Pe$, and to the stationary instability on decreasing $\beta$. Note that the dependence of the dynamics on the asymmetry $\beta$ can be rationalized as follows: a negative determinant indicates that the signs of the two eigenvalues are different, and so a negative $\Delta(J)$ cannot yield imaginary eigenvalues (as they necessarily have to be complex conjugates as the Jacobian matrix is real). This implies the result that for small $\alpha$, $\beta>1$ can yield $\Delta(J)>0$, and so there can be modes that have oscillatory instability only for $\beta>1$.

### v2

# From Order to Chimeras: Unraveling Dynamic Patterns in Active Fluids with Nonlinear Growth

@@ 1 Introduction
Pattern formation is a ubiquitous phenomenon in natural systems. In particular, it is an integral part of the development of biological systems. The classical framework for modelling and understanding pattern formation is that of reaction-diffusion systems [1, 2, 3, 4, 5]. The emergent patterns observed in this broad class of systems include time-independent and time-dependent oscillatory patterns [1], symmetry-breaking instabilities [6], travelling waves [7], spirals [8] and jumping oscillations [9]. Importantly, many of these patterns have been verified in experiments as well [10, 11].
However, in most biological systems, active transport and forces play a critical role in pattern formation. A prototypical approach to understanding pattern formation in biological systems, therefore, integrates the contributions of mechanical and chemical effects [12], for instance actin networks showing non-equilibrium dynamics by force generation through myosin motor activity [13, 14, 15] and pattern formation in active fluid medium in the presence of diffusing chemical species that are advected by self-generated flows produced by concentration-dependent active stress gradients [16, 13]. Cell crawling is also dependent on the density of myosin motors, and so it can be studied by considering the distribution of myosin motors as a supercritical van der Waals (vdW) fluid [17]. Spontaneous protrusion dynamics is modelled by mechano-chemical coupling via a polymerizing active gel layer [18, 19].
@@
The actomyosin cortex, which lies just beneath the cell membrane, consists of actin filaments and myosin motor proteins which crosslink the actin. This generates mechanical forces, giving rise to an active stress component in a thin mesoscopic layer of the actomyosin cortex. The actomyosin cortex is treated as an active fluid at time scales of morphogenesis, which is an essential part of organismal development, relevant to human tissue development as well as development in systems such as worms, flies, zebrafish and mice [20]. The active stress can now be considered to be a function of the concentrations of regulatory chemical species to complete the mechanochemical integration, leading to pattern formation even in the absence of chemical reactions [16]. When extended to two chemical species which undergo advection-diffusion and regulate the active stress, this framework gives rise to pulsatory patterns [13]. Spontaneously emerging localized states in the active fluid medium have led to the understanding of localized cellular patterns [21]. It has been shown that localized states can emerge spontaneously (analogous to isolated clusters of actin and signalling molecules in cancer cells [22]) if the assembly of active matter is regulated by the presence of chemical species that are advected with flows resulting from active stress gradients.
@@
In another research direction, in the context of dynamical systems in general, a class of patterns that have generated intense research interest are chimera states. The scope of the term “chimera” has vastly expanded since it was first introduced [23], and a number of variations such as frequency chimeras, amplitude and amplitude mediated chimeras, breathing chimeras, traveling phase-cluster chimeras and multiple phase-cluster chimeras, have been characterized. These states have had alternate descriptors in earlier papers such as clustered states, or coexisting states. Broadly, a chimera state refers to one where a system spontaneously breaks the underlying symmetry and splits into co-existing groups that have significantly different dynamical features [24, 25, 26], i.e. the emergent pattern has spatial domains exhibiting distinct temporal behaviour. The term “chimera-like” has also been widely used to refer to non-stationary chimera states with irregular phase boundaries [27, 28]. This fascinating phenomenon has been observed in a variety of systems, ranging from Josephson junction arrays [29] to uni-hemispheric sleep in certain animals [30], as well as in continuous media models [31, 32]. Importantly, these chimera states have also been observed experimentally in optical analogs of coupled map lattices [33], Belousov-Zhabotinsky chemical oscillator systems [34], two populations of mechanical metronomes [35], electronic circuit systems [36, 37], neurodynamics [38] and Liquid Crystal Light Valve experiments with optical feedback [39]. So, it is of immense interest to uncover chimera patterns in different classes of systems to gauge the generality of this interesting spatiotemporal pattern.
Motivated by these broad ideas, in this paper, we show the emergence of chimera states in an active fluid with two chemical species regulating the active stress as in Ref. [13], with one species showing a logistic growth. Our work, while generic for any two-component advection-diffusion system coupled to an active fluid, is more specifically related to the actomyosin cortex. Pulsatile patterns have been observed in several situations in the actomyosin cytoskeleton. In that context, myosin motor proteins could represent the species that up-regulates stress. Any of the several proteins associated with the actomyosin cortex - like actin-monomer binding proteins, severing proteins, cross-linking proteins or filament binding proteins - could serve as the species which down-regulates the stress. The logistic growth term also has a simple explanation. In pattern formation in actin networks, the growth term arises naturally to describe the growth of existing F-actin fibres and nucleation of new fibres, while crowding will slow down polymerization to justify the satuaration [40]. Separately, a generic mechanism involving actin turnover and myosin activity in the ring formation in the Drosophila trachea also necessitates using a variant of the non-linear growth term [41]. Both these settings fall within the necessary requirements in our model.
@@
A logistic growth term has also been used extensively in many mathematical models of biological systems and provides a general description of population growth. In particular, a logistic growth model with long-range interactions serves as a generic minimal model for competition for common resources and pattern formation in excitable media [42]. For instance, in the chemotaxis model proposed in Ref. [43], the logistic term is used to model the growth of cell density. In autochemotactic pattern formation, logistic growth models the death and reproduction of self-propelled bacteria  [44]. Further, the logistic functional form has been used in an experimentally motivated model of the coupling between reaction-diffusion and active matter [45], as well as in gene expression dynamics [46, 47]. Polymerization in actin gels and solutions is also expressed by a reactive logistic term, useful to describe many characteristic states of actin-wave formation: spots, spirals, and travelling waves [48, 40].
@@ 2 Mathematical Model
Thus, both species undergo advection and diffusion, with the diffusive component determined by the diffusion coefficient $D$ and the advective component given by the bulk fluid flow velocity $v$. The relative diffusion coefficient of the two species is determined by the parameter $\alpha>0$. The nonlinear logistic growth term, with strength $r>0$, has the capacity to destabilize a steady state with zero concentration and also yields saturation at a finite carrying capacity $K$. For the active fluid at low Reynolds numbers, the inertial terms can be neglected, and the force balance equation gives
@@
where $\beta$ is an asymmetry parameter, and $A_{S}$ and $I_{S}$ represent the saturation values of the concentrations of the two chemical species [13]. For $\beta<-1$, $A$ down-regulate and $I$ up-regulates active stress; for $1\leq\beta\leq 1$ both species up-regulates stress and for $\beta>1$, $A$ up-regulates while $I$ down-regulates active stress. For $A<A_{S}$ and $I<I_{S}$, this functional form is linear in leading order, and a simple linear functional form of $f(A,I)$ also exhibits qualitatively similar spatiotemporal patterns.
@@
$\partial_{t}A=-\partial_{x}(vA)+\partial^{2}_{x}A+RA(1-A)$ (7)
@@
3 Stability Analysis
Figure 1: (a) Phase diagram of the different dynamical patterns that emerge, in the parameter space of the Péclet number $Pe$ and the scaled nonlinear growth parameter $R$, obtained from Linear Stability Analysis. Here the ratio of diffusion constants $\alpha=0.1$ and the asymmetry parameter $\beta=3$. The brown colour represents the stable homogeneous steady state, pale blue represents oscillatory instability, and yellow represents stationary instability. The black curve represents the linear boundary between the stable homogeneous steady state and the regime of oscillatory instability, given by Eqn. 13. Further, the best-fit linear boundary, $Pe\sim 34R$, between the regime of oscillatory instability and stationary instability is also demarcated in black. (Inset) Numerically obtained bifurcation diagram, with PePe varying from $11$ to $14$ ($\beta$=33, $\alpha$=$0.1$, $R=1$), clearly showing the transition from homogeneous steady states to oscillatory regime, through a supercritical Hopf bifurcation. Here, $A_{m}$ represents the local extrema of $A$ at a particular site after a long transience. (b-d) Kymographs obtained from numerical simulations, illustrating the emergent spatiotemporal patterns in the region of instability, for $R=1$, as the Péclet number increases from left to right: (b) $Pe=18$ (regular pulsatory state), (c) $Pe=35$ (chimera state) and (d) $Pe=48$ (spatiotemporal chaos). In the kymographs, the concentration $A$ is displayed as a function of time (vertical axis) and space (horizontal axis), with the colour bar indicating increasing concentrations from $A=0$ (black) to $A=2$ (white).
Consider homogeneous steady states, with concentrations $(A_{0},I_{0})$, and velocity $v=0$. We consider perturbations to the homogeneous steady state $(A_{0},I_{0})$ of the form $A=A_{0}+A_{p}(x,t)$, $I=I_{0}+I_{p}(x,t)$ and $v=v_{0}+v_{p}(x,t)$, where, $A_{p}(x,t)=\delta A_{0}e^{\lambda t+ikx}$, $I_{p}(x,t)=\delta I_{0}e^{\lambda t+ikx}$, $v_{p}(x,t)=\delta v_{0}e^{\lambda t+ikx}$. The important parameters here are (1) the Péclet number, which gives the ratio of the diffusive time scale and the advective time scale, (2) the parameter $R$, which reflects the strength of nonlinearity, and (3) $\beta$, which reflects the asymmetry in the two chemical species and influences evolution through the function $f$. Linear stability analysis of Eqs. (7)-(9), yields the following Jacobian:
@@ 3 Stability Analysis
where $f_{A},f_{I}$ are the partial derivatives of $f$ with respect to $A$ and $I$ respectively, evaluated at the homogeneous steady state. In what follows we consider $(A_{S},I_{S})=(3A_{0},3I_{0})$ in accordance with the choice in Ref. [13]. Note that the qualitative nature of the emergent patterns is not sensitive to this choice.
For each value of wave number $k=\frac{2\pi n}{L}$, $n=0,1,\dots$ (considering periodic boundary conditions), we obtain two eigenvalues, which determine the stability and nature of the emergent dynamical pattern. The trace and the determinant of the Jacobian are given as:
@@
Now, for a homogeneous steady state to be stable under perturbation, the disturbances from the homogeneous profile must damp exponentially rapidly in time. So the real parts of the eigenvalues for all $k$ should be less than zero, implying that $Tr(J)<0$ and $\Delta(J)>0$. In the parameter regime where the uniform steady state is unstable, the real part of at least one of the eigenvalues is positive, leading to the formation of diverse patterns. We denote the regime where the leading eigenvalues may not be imaginary as a stationary instability, in analogy with the terminology used in the Turing model. If the eigenvalues are imaginary for some values of $k$, an oscillatory instability will emerge.
Figure 2: (a) Phase diagram of the different dynamical patterns that emerge, in the parameter space of the Péclet number $Pe$ and asymmetry parameter $\beta$, obtained from Linear Stability Analysis, for $\alpha$=$0.1$ and scaled nonlinear growth parameter $R=1$. The brown colour represents the stable homogeneous steady state, pale blue represents oscillatory instability, and yellow represents stationary instability. Symbols in the phase diagram denote the dynamical patterns obtained through numerical simulations: circle represents homogeneous steady states, square represents stationary Turing-like patterns, triangle represents pulsatory patterns, diamond represents chimera states and star represents merging-emerging dynamics. (b-d) Kymographs obtained from numerical simulations illustrating the emergent spatiotemporal patterns in the region of instability: (b) $\beta=0.5,Pe=5$ (stationary patterns, denoted by squares in the phase diagram); (c) $\beta=1.75$, $Pe=14$ (pulsatory pattern, denoted by triangles in the phase diagram); (d) $\beta=1.75$, $Pe=17$.
@@
We generate phase diagrams in the space of different parameter sets (cf Figs 1-2) considering the variation of the eigenvalues with respect to wavenumber $k$, and noting the sign of the real part of the leading eigenvalue (see dispersion curves in Fig. 3 which illustrate the behaviour of the real and imaginary parts of the leading eigenvalue when a parameter crosses a critical value giving rise to different classes of transitions). The three distinct dynamical phases, namely the stable homogeneous steady state, the regime of oscillatory instability and the regime of stationary instability, are marked with different colors in the phase diagrams.
@@
In Fig. 1, we display the phase diagram in the parameter space of the Péclet number $Pe$ and the scaled nonlinear growth parameter $R$. It is apparent from the phase diagram that as the Péclet number increases beyond a critical value, the homogeneous steady state gets destabilized, giving way to oscillatory instability. Inspecting the behaviour of the leading eigenvalue as the Péclet number crosses the critical value reveals a supercritical transition from the stable homogeneous steady state to oscillatory instability. On further increase of the Péclet number, the oscillatory instability changes to a stationary instability.
@@
In contrast, interestingly, the phase diagram also shows that the parameter space occupied by the homogeneous steady state is enlarged under increasing $R$, i.e. larger nonlinear growth yields enhanced homogeneity. This indicates a counter-intuitive stabilizing effect of increasing non-linearity on the dynamics. Further, increasing the magnitude of the nonlinear growth term also increases the parameter region supporting oscillatory states. It is also distinctly evident from Figure 1 that the boundaries between the different dynamical regimes have an approximate linear dependence on the scaled nonlinearity $R$. The slope of the linear rise of the boundary curve separating the homogeneous steady state and the oscillatory state is gentler than that of the boundary between the stationary instability and oscillatory instability. This implies that the oscillatory state increases most significantly with increasing nonlinear growth.
@@
Specifically, our analysis also yields the boundary curve between the stable homogeneous steady state and the oscillatory instability. Along this boundary curve, the condition $Tr(J)=0$ should be satisfied. So from the expression in Eqn. 11, we have $Tr(J)=-k^{2}(1+\alpha)-R+\frac{Pek^{2}(f_{A}+f_{I})}{(1+k^{2})}=0$, i.e. $k^{2}(1+\alpha)+R=\frac{Pek^{2}(f_{A}+f_{I})}{(1+k^{2})}$. Now, as in the phase diagram in Fig. 1, we consider the values of $\alpha=0.1$ and $\beta=3$ (which gives $f_{A}+f_{I}=\frac{3}{8}$). Further, taking $k=1$ (since $k_{max}\sim 1$ over a large range of Péclet numbers) we obtain: $(1+\alpha)+R=\frac{3Pe}{16}$. This yields the functional relation between the critical $Pe$ number and $R$ to be:
@@
This linear relation is indicated by a line in the phase diagram displayed in Fig. 1, and it is clear that this approximate expression fits results from stability analysis over a large range of $R$. Since the critical value of the Péclet number increases almost linearly with the scaled nonlinear growth parameter $R$, it implies that the range of $Pe$ over which the homogeneous steady state is stable is directly proportional to $R$, with larger stable ranges emerging for larger $R$.
When the system is weakly nonlinear, the critical value of the $Pe$ number also becomes very small, with $Pe_{c}=5.87$ for $R\to 0$, from Eqn. 13. However, notice that there is a small deviation from the linear relation given in Eqn. 13, in the phase boundary in Fig. 1, in the limit of weak nonlinearity. This can be rationalized as follows: Eqn. 13 was obtained using the approximation that $k_{max}\sim 1$. However, for very low $Pe$, this approximation is not accurate. Using the exact explicit expression for $k_{max}$ in place of the approximation yields a nonlinear equation for the critical $Pe$ number. Solving this nonlinear equation numerically (or graphically) gives a more accurate estimate of the critical $Pe$ number for weak nonlinearity, with $Pe_{c}=8(1+\alpha)/3=8.8/3\sim 3$ for $R\to 0$. This is completely consistent with the phase diagram in Fig. 1. For higher nonlinearity $R$, $Pe_{c}$ is also larger, and the linear relation in Eqn. 13 then offers an excellent approximation to the behaviour of the stability boundary.
@@
Figure 3: Dispersion curves obtained from Linear Stability Analysis, for varying Péclet number $Pe$ and asymmetry parameter $\beta$, illustrating the transitions between the different dynamical regimes. Here $\alpha=0.1$ and nonlinearity parameter $R=1$. The top row of panels show the transition from the stable homogeneous steady state to the oscillatory instability, as PePe increases from 99 to $15$, for $\beta=1.2$. The middle row of panels shows the transition from the stable homogeneous steady state to stationary instability as $\beta$ decreases from $1.1$ to $0.7$ for fixed Péclet number $Pe=6$. The bottom row of panels shows the transition from oscillatory instability to stationary instability, as PePe increases from $13$ to $20$, for $\beta=1.2$. In all panels, the blue curve represents the real part of the leading eigenvalue, the red curve represents the imaginary part of the leading eigenvalue, and the black dotted line shows $\lambda=0$ for reference.
@@
We also observe that the pattern formation is crucially dependent on the asymmetry parameter $\beta$. This is clearly evident from Figure 2, where we display phase diagrams in the parameter space of the Péclet number $Pe$ and asymmetry $\beta$ for different values of the scaled logistic growth parameter $R$. When $\beta$ is low ($\beta<1$), no regime of oscillatory instability exists for any value of the Péclet number, while for $\beta$ greater than $1$, all three dynamical regimes are possible. Another interesting observation is the independence of the phase boundary on the asymmetry parameter $\beta$, after $\beta\sim 1$ in Fig. 2. This can be rationalized as follows: the parameter $\beta$ enters the expression for the Jacobian through $f_{A}$ and $f_{I}$. Now the Trace is actually independent of $\beta$, as $(f_{A}+f_{I})=3/8$, making the expression in Eq. 11 independent of $\beta$. The Determinant, however, involves $\beta$. When $\beta<1$, $f_{I}=\frac{3}{16}(1-\beta)$ is positive, and so the stability boundary is dependent on $\beta$ through the sign of the Determinant which is determined by the relative magnitudes of the positive and negative terms in Eqn. 12. On the other hand, for $\beta>1$, $f_{I}$ is always negative, and so the sign of the Determinant is always positive after $\beta\sim 1$, in the limit of small $\alpha$. This implies that after $\beta\sim 1$, $\beta$ no longer influences the eigenvalues, and so the critical $Pe$ number is independent of it.
Further, sets of dispersion curves are also displayed in Fig. 3, clearly showing the behaviour of the real and imaginary parts of the leading eigenvalues when $Pe$ and $\beta$ cross critical values, underlying different classes of dynamical transitions. In particular, these dispersion curves illustrate the transitions from the stable homogeneous steady state to oscillatory instability on increasing $Pe$, and to the stationary instability on decreasing $\beta$. Note that the dependence of the dynamics on the asymmetry $\beta$ can be rationalized as follows: a negative determinant indicates that the signs of the two eigenvalues are different, and so a negative $\Delta(J)$ cannot yield imaginary eigenvalues (as they necessarily have to be complex conjugates as the Jacobian matrix is real). This implies the result that for small $\alpha$, $\beta>1$ can yield $\Delta(J)>0$, and so there can be modes that have oscillatory instability only for $\beta>1$.